OpenStudy (andijo76):

Given that P(A) = 0.5, P(B) = 0.3, and P(A and B) = 0.10, determine P(A|B) 0.125 0.20 0.33 or 0.50

OpenStudy (anonymous):

\[P(A|B)=\frac{P(A\cap B)}{P( B)}\] plug in the numbers and divide

OpenStudy (anonymous):

\[\frac{.1}{.3}=\frac{1}{3}\]